OpenStudy (anonymous):

Which are solutions to y=x-4? Choose all correct answers. (4,0), (3,-1), (6,3), (2,-4)

OpenStudy (anonymous):

To answer such questions you should plug in. you plug in by replacing (x,y) with the values in the brackets, so y=x-4; let's try the solutions A; 0=4-4=0 ; so A works B; -1=3-4=-1 so B works c; 3=6-4=2 which is false, so C doesn't work D; -4=2-4=-2 which is false so D doesn't work. Ciao!

OpenStudy (anonymous):

Thank you :)